OpenStudy (haichi):

Alfred is taking a trip. He drove 150 mi and then stopped to refuel. He will drive 60 mph to finish his trip. Let x represent the hours it takes for Alfred to complete the rest of his trip, and let y represent the total distance traveled. Complete the function equation that relates Alfred’s total distance traveled to his time.

OpenStudy (haichi):

I donnt understand

OpenStudy (haichi):

Dont*

OpenStudy (haichi):

@jim_thompson5910

OpenStudy (haichi):

@Nnesha

OpenStudy (haichi):

thnx @jagr2713

jagr2713 (jagr2713):

yw

jimthompson5910 (jim_thompson5910):

He already drove 150 miles He'll drive another additional 60x miles. The 60 is from the 60 mph and the x is the number of hours. So (speed)*(time) = 60*x = 60x is the distance he drives after stopping for gas.

jimthompson5910 (jim_thompson5910):

so you just add the two to get the total distance y =150 + 60x y = 60x + 150

OpenStudy (haichi):

awsome

OpenStudy (haichi):

btw do u have ur own tutoring site?

OpenStudy (haichi):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Yes I do

OpenStudy (haichi):

Got the link?

jimthompson5910 (jim_thompson5910):

http://www.freewebs.com/jimthompson5910/home.html

OpenStudy (haichi):

thnks u

jimthompson5910 (jim_thompson5910):

np